Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
kentaro.komori - 22:38 Wednesday 23 August 2023 (26455) Print this report
Putting additional RFPDs on the REFL optics table

[Ishikawa, Tanaka, Komori]

We placed additional RFPDs on the REFL table, but we were not able to complete the task.
The remaining tasks will be finished tomorrow morning.

What We Accomplished:

  • Installed all the additional optical components on the table.
  • Established connections with the corresponding ports (klog:26361) at the rack side.
  • Verified the functionality of the DC signals from the two RFPDs.

Plan for Tomorrow:

  • Thread the remaining cables through the hole for signal cables and establish connections at the table side.
  • Perform power tuning and alignment.
Comments to this report:
kentaro.komori - 22:40 Thursday 24 August 2023 (26474) Print this report

Today, we successfully completed this task.
We dump the ghost beams originating from the surfaces of the two RFPDs.
Additionally, we isolated the PD case ground from the optics table by using vinyl tape.

The input powers for RFPD3 (CARM out of loop) and RFPD4 (MICH) were measured under the condition of 0.8 W IMC output.
The results are as follows:

  • RFPD3: 3.8 mW
  • RFPD4: 6.3 mW

The variation arises due to the Thorlab beam splitter imbalance for p-polarized light.

Upon increasing the IMC output to 4 W, the input powers for the PDs are anticipated to be approximately 20 mW and 30 mW, respectively.
These powers will drop by one-third when the interferometer is locked, resulting in observation power levels of around 7 mW and 10 mW.

If required, the power on the PDs can be increased by adjusting the Half-Wave Plate (HWP) angle.

I will provide an updated layout for the REFL optical layout shortly.

Search Help
×

Warning

×